# Hierarchical Selective Classification

### v1

@@ Abstract
Deploying deep neural networks for risk-sensitive tasks necessitates an uncertainty estimation mechanism. This paper introduces hierarchical selective classification, extending selective classification to a hierarchical setting. Our approach leverages the inherent structure of class relationships, enabling models to reduce the specificity of their predictions when faced with uncertainty. In this paper, we first formalize hierarchical risk and coverage, and introduce hierarchical risk-coverage curves. Next, we develop algorithms for hierarchical selective classification (which we refer to as “inference rules”), and propose an efficient algorithm that guarantees a target accuracy constraint with high probability. Lastly, we conduct extensive empirical studies on over a thousand ImageNet classifiers, revealing that training regimes such as CLIP, pretraining on ImageNet21k and knowledge distillation boost hierarchical selective performance.
@@ 1 Introduction
Deep neural networks (DNNs) have achieved incredible success across various domains, including computer vision and natural language processing. To ensure the reliability of models intended for real-world applications we must incorporate an uncertainty estimation mechanism, such as selective classification [17], which allows a model to abstain from classifying samples when it is uncertain about their predictions. Standard selective classification, however, has a shortcoming: for any given sample, it is limited to either generating a prediction or rejecting it, thereby ignoring potentially useful information about the sample, in case of rejection.
To illustrate the potential consequences of this limitation, consider a model trained for classifying different types of brain tumors from MRI scans, including both benign and malignant tumors. If a malignant tumor is identified with high confidence, immediate action is needed. For a particular hypothetical scan, suppose the model struggles to distinguish between 3 subtypes of malignant tumors, assigning a confidence score of 0.33 to each of them (assuming those estimates are well-calibrated and sum to 1). In the traditional selective classification framework, if the preset confidence threshold is higher than 0.33, the model will reject the sample, failing to provide valuable information to alert healthcare professionals about the patient’s condition, even though the model has enough information to conclude that the tumor is malignant with high certainty. This could potentially result in delayed diagnosis and treatment, posing a significant risk to the patient’s life. However, a hierarchically-aware selective model with an identical confidence threshold would classify the tumor as malignant with 99% confidence. Although this prediction is less specific, it remains highly valuable as it can promptly notify healthcare professionals, leading to early diagnosis and life-saving treatment for the patient.
This motivates us to propose hierarchical selective classification (HSC), an extension of selective classification to a setting where the classes are organized in a hierarchical structure. Such hierarchies are typically represented by a tree-like structure, where each node represents a class, and the edges reflect a semantic relationship between the classes, most commonly an ’is-a’ relationship. Datasets with an existing hierarchy are fairly common, for instance, the ImageNet dataset [10], which is widely used in computer vision tasks, is organized according to the WordNet hierarchy [29]. A visualization of a small portion of the ImageNet hierarchy is shown in Figure 1.
@@
(4) We conduct a comprehensive empirical study evaluating HSC on more than 1,000 ImageNet classifiers. We present numerous previously unknown observations, most notably that training approaches such as pretraining on larger datasets, contrastive language-image pretraining (CLIP) [31], and knowledge distillation significantly boosts selective hierarchical performance.
@@ 2 Problem Setup
Selective Classification Let $\mathcal{X}$ be the input space, and $\mathcal{Y}$ be the label space. Samples are drawn from an unknown joint distribution $P(\mathcal{X}\times\mathcal{Y})$ over $\mathcal{X}\times\mathcal{Y}$. A classifier $f$ is a prediction function $f:\mathcal{X}\rightarrow\mathcal{Y}$, and $\hat{y}=f(x)$ is the model’s prediction for $x$. The true risk of a classifier $f$ w.r.t. $P$ is defined as: $R(f|P)=\mathbb{E}_{P}[\ell(f(X),Y)]$, where $\ell:\mathcal{Y}\times\mathcal{Y}\rightarrow\mathbb{R}^{+}$ is a given loss function, for instance, the 0/1 loss. Given a set of labeled samples $S_{m}=\{(x_{i},y_{i})\}_{i=1}^{m}$, the empirical risk of $f$ is: $\hat{R}(f|S_{m})=\frac{1}{m}\sum_{i=1}^{m}\ell(f(x_{i}),y_{i}).$ Following the notation of [19], we use a confidence score function $\kappa(x,\hat{y}|f)$ to quantify prediction confidence. We require $\kappa$ to induce a partial order over instances in $\mathcal{X}$. In this work, we focus on the most common and well-known $\kappa$ function, softmax response [9, 35]. For a classifier $f$ with softmax as its last layer: $\kappa(x,\hat{y}|f)=f_{\hat{y}}(x)$. Softmax response has been shown to be a reliable confidence score in the context of selective classification [16, 17] as well as in hierarchical classification [40], consistently achieving good performance.
@@
Hierarchical Classification Following the notations of [11] and [40], a hierarchy $H=(V,E)$ is defined by a tree, with nodes $V$ and edges $E$, the root of the tree is denoted $r\in V$. Each node $v\in V$ represents a semantic class: the leaf nodes $\mathcal{L}\subseteq V$ are mutually exclusive ground-truth classes, and the internal nodes are unions of leaf nodes determined by the hierarchy. The root represents the semantic class containing all other objects. A sample $x$ that belongs to class $y$ also belongs to the ancestors of $y$. Each node has exactly one parent and one unique path to it from the root node. The set of leaf descendants of node $v$ is denoted by $\mathcal{L}(v)$, and the set of ancestors of node $v$, including $v$ itself, is denoted by $\mathcal{A}(v)$. A hierarchical classifier $f:\mathcal{X}\rightarrow V$ labels a sample $x\in\mathcal{X}$ as a node $v\in V$, at any level of the hierarchy, as opposed to a flat classifier, that only predicts leaves.
@@
Hierarchical Selective Classification Selective classification offers a binary choice: either to predict or completely reject a sample. We propose hierarchical selective classification (HSC), a hierarchical extension of selective classification, which allows the model to retreat to less specific nodes in the hierarchy in case of uncertainty. Instead of rejecting the whole prediction, the model can now partially reject a sample, and the degree of the rejection is determined by the model’s confidence. For example, if the model is uncertain about the specific dog breed of an image of a Labrador but is confident enough to determine that it is a dog, the safest choice by a classic selective framework would be to reject it. In our setting, however, the model can still provide useful information that the object in the image is a dog (see Figure 1).
@@
To our knowledge, controlling the trade-off between accuracy and specificity was only previously explored by [11]. They proposed the Dual Accuracy Reward Trade-off Search (DARTS) algorithm, which attempted to obtain the most specific classifier for a user-specified accuracy constraint. They used information gain and hierarchical accuracy as two competing factors, integrating them into a generalized Lagrange function. Our approach differs from theirs in that we guarantee control over the full trade-off, while some coverages cannot be achieved by DARTS.
As mentioned in section 2, it is considered correct to label an image as either its ground truth leaf node or any of its ancestors. Thus, we employ a natural extension of the 0/1 loss to define the true risk of a hierarchical classifier $f^{H}$ with regard to $P$: $R^{H}(f^{H}|P)=\mathds{E}_{P}[f^{H}(X)\notin\mathcal{A}(Y)],$ and the empirical risk over a labeled set of samples $S_{m}$: $\hat{R^{H}}(f^{H}|S_{m})=\frac{1}{m}\sum_{i=1}^{m}\mathds{1}[f^{H}(x_{i})\notin\mathcal{A}(y_{i})].$ To ensure that specific labels are preferred, it is necessary to consider the specificity of predictions. Hierarchical selective coverage, which we propose as a hierarchical extension of selective coverage, measures the amount of information present in the model’s predictions. A natural quantity for that is entropy, which measures the uncertainty associated with the classes beneath a given node. Assuming a uniform prior on the leaf nodes, the entropy of a node $v\in V$ is $H(v)=-\sum_{{v^{\prime}\in\mathcal{L}(v)}}\frac{1}{|\mathcal{L}(v)|}\log(\frac{1}{|\mathcal{L}(v)|})=\log(|\mathcal{L}(v)|).$ At the root, the entropy reaches its maximum value, $H(r)=\log(|\mathcal{L}|)$, while at the leaves the entropy is minimized, with $H(y)=0$ for any leaf node $y\in\mathcal{L}$. This allows us to define coverage for a single node $v$, regardless of $g_{\theta}^{H}$. We define hierarchical coverage (from now on referred to as coverage) as the entropy of $v$ relative to the entropy of the root node: ${\phi}^{H}(v)=1-\frac{H(v)}{H(r)}.$ The root node has zero coverage, as it does not contain any information. The coverage gradually increases until it reaches 1 at the leaves. We can also define true coverage for a selective model: $\phi^{H}(f^{H})=\mathds{E}_{P}[\phi(f^{H}(X)].$ The empirical coverage of a classifier over a labeled set of samples $S_{m}$ is defined as the mean coverage over its predictions: $\hat{\phi^{H}}(f^{H}|S_{m})=\frac{1}{m}\sum_{i=1}^{m}\phi(f^{H}(x_{i})).$ For a hierarchy comprised of leaves and a root node, hierarchical coverage reduces to selective coverage, where classifying a sample as the root corresponds with rejection.
@@ 3 Hierarchical Selective Inference Rules
To define a hierarchical selective model $f^{H}=(f,g_{\theta}^{H})$ given a base classifier $f$, an explicit definition of $g_{\theta}^{H}$ is required. $g_{\theta}^{H}$ determines the internal node in the hierarchy $f^{H}$ retreats to when the leaf prediction of $f$ is uncertain. $g_{\theta}^{H}$ requires obtaining confidence for internal nodes in the hierarchy. Since most modern classification models only assign probabilities to leaf nodes, we follow [11] by setting the probability of an internal node to be the sum of its leaf descendant probabilities: $f^{H}_{v}(x)=\sum_{y\in\mathcal{L}(v)}f_{y}(x).$ Unlike other works that assign the sum of leaf descendant probabilities to internal nodes, our algorithm first calibrate the leaf probabilities through temperature scaling. Since the probabilities of internal nodes heavily rely on the leaf probabilities supplied by $f$, calibration is beneficial for obtaining more reliable leaf probabilities, and since the internal nodes probabilities are sums of leaf probabilities, this cumulative effect is even more pronounced. Further, [16] found that applying temperature scaling also improves ranking and selective performance, which is beneficial to our objective. In this section, we introduce several inference rules, along with useful theoretical properties. We propose the Climbing inference rule (Algorithm 1), which starts at the most likely leaf and climbs the path to the root, until reaching an ancestor with confidence above the threshold. A visualization of the Climbing inference rule is shown in Figure 1.
We compare our proposed inference rules to the following baselines: The first is the selective inference rule, which predicts ’root’ for every uncertain sample (this approach is identical to standard “hierarchically-ignorant” selective classification). The second hierarchical baseline, proposed by [40], selects the node with the highest coverage among those with confidence above the threshold. We refer to this rule as “Max-Coverage” (MC), and its algorithm is detailed in Appendix A.
@@
The Climbing inference rule outlined in Algorithm 1 satisfies both monotonicity properties. MC satisfies monotonicity in coverage, but not in correctness. An additional algorithm that does not satisfy neither of the properties is discussed in Appendix B.
@@ 4 Optimal Selective Threshold Algorithm
3:Based on $\alpha,n,\delta$, calculate $\epsilon$triangleright See Appendix D
@@
6:     $\theta_{i}\leftarrow(g_{\theta}^{H})^{-1}(x_{i}|\kappa,f)$triangleright See Algorithm 5 in Appendix C
@@
Proof: See Appendix D.
Note: our algorithm can provide an even greater degree of flexibility: the user may choose to set the values of any three parameters out of $\alpha,n,\epsilon,\delta$. With the three parameters fixed, we can compute the remaining parameter. In the algorithm we implicitly assumed $n$ is fixed, but increasing $n$ yields more stable results. For a detailed discussion of this topic, see Appendix D.
5 Experiments
In this section we evaluate the methods introduced in Section 3 and Section 4. The evaluation was performed on 1115 vision models pretrained on ImageNet1k [10], and 6 models pretrained on iNat-21 [22] (available in timm 0.9.16 [43] and torchvision 0.15.1 [28]). The reported results were obtained on the corresponding validation sets (ImageNet1k and iNat-21). The complete results and source code necessary for reproducing the experiments are provided in the Supplementary Material.
Inference Rules: Table 1 compares the mean results of the Climbing inference rule to both hierarchical and non-hierarchical baselines. The evaluation is based on RC curves generated for 1115 models pretrained on ImageNet1k and 6 models pretrained on iNat21 with each inference rule applied to their output. These aggregated results show the effect of different hierarchical selective inference rules, independent of the properties or output of a specific model. Compared to the non-hierarchical selective baseline, hierarchical inference has a clear benefit. Allowing models to handle uncertainty by partially rejecting a prediction instead of rejecting it as a whole, proves to be advantageous. The average model is capable of leveraging the hierarchy to predict internal nodes that reduce risk while preserving coverage. Nonetheless, the differences remain stark when comparing the hierarchical inference rules. Climbing, achieving almost 15% hierarchical gain, outperforms MC, with more than double the gain of the latter. These results highlight the fact that the approach taken by inference rules to leverage the hierarchy can significantly impact outcomes. A possible explanation for Climbing’s superior performance could stem from the fact that most models are trained to optimize leaf classification. By starting from the most likely leaf, Climbing utilizes the prior knowledge embedded in models for leaf classification, while MC ignores it. Interestingly, our analysis also shows that HSC improves confidence calibration. Appendix F discusses the calibration of HSC in more detail.
Table 1: Comparison of mean hAURC and hierarchical gain results for the selective, Max-Coverage (MC) and Climbing inference rules applied to 1115 models trained on ImageNet1k, and 6 models trained on iNat21.
ImageNet-1k (1115 models) iNat21 (6 models)
@@ 5 Experiments
Optimal Threshold Algorithm: We compare our algorithm to DARTS [11]. We evaluate the performance of both algorithms on 1115 vision models trained on ImageNet1k, for a set of target accuracies. For each model and target accuracy the algorithm was run 1000 times, each with a randomly drawn calibration set. We also evaluate both algorithms on 6 models trained on the iNat21 dataset. We compare the algorithms based on two metrics: (1) target accuracy error, i.e., the mean distance between the target accuracy and the accuracy measured on the test set; (2) coverage achieved by using the algorithms’ output on the test set. The results presented in Table 2, and Table 4 in Appendix G, show that our algorithm consistently achieves substantially lower target accuracy error, indicating that our algorithm succeeds in nearing the target accuracy more precisely. This property allows the model to provide a better balance between risk and coverage. Our algorithm is more inclined towards this trade-off, as it almost always achieves higher coverage than DARTS. This is particularly noteworthy when the desired target is high: while DARTS loses coverage quickly, our algorithm manages to maintain coverage that is up to twice as high. Importantly, DARTS does not capture the whole risk-coverage trade-off. Specifically, at the extreme point in the trade-off when it aims to maximize specificity, it still falls short of providing full coverage, and its predictions do not reduce to a flat classifier’s predictions, i.e. it may still predict internal nodes.

### v2

@@ Abstract
Deploying deep neural networks for risk-sensitive tasks necessitates an uncertainty estimation mechanism. This paper introduces hierarchical selective classification (HSC), extending selective classification to a hierarchical setting. Our approach leverages the inherent structure of class relationships, enabling models to reduce the specificity of their predictions when faced with uncertainty. In this paper, we first formalize hierarchical risk and coverage, and introduce hierarchical risk-coverage curves. Next, we develop algorithms for performing HSC (which we refer to as “inference rules”), and propose an efficient algorithm that guarantees a target accuracy constraint with high probability. We also introduce Calibration-Coverage curves, which analyze the effect of hierarchical selective classification on calibration. Our extensive empirical studies on over a thousand ImageNet classifiers reveal that training regimes such as CLIP, pretraining on ImageNet21k, and knowledge distillation boost hierarchical selective performance. Lastly, we show that HSC improves both selective performance and confidence calibration. Code is available at https://github.com/shanigoren/Hierarchical-Selective-Classification.
@@ 1 Introduction
Deep neural networks (DNNs) have achieved incredible success across various domains, including computer vision and natural language processing. To ensure the reliability of models intended for real-world applications we must incorporate an uncertainty estimation mechanism, such as selective classification [19], which allows a model to abstain from classifying samples when it is uncertain about their predictions. Standard selective classification, however, has an inherent shortcoming: for any given sample, it is limited to either making a prediction or rejecting it, thereby ignoring potentially useful information about the sample, in case of rejection.
To illustrate the potential consequences of this limitation, consider a model trained for classifying different types of brain tumors from MRI scans, including both benign and malignant tumors. If a malignant tumor is identified with high confidence, immediate action is needed. For a particular hypothetical scan, suppose the model struggles to distinguish between 3 subtypes of malignant tumors, assigning a confidence score of 0.33 to each of them (assuming those estimates are well-calibrated and sum to 1). In the traditional selective classification framework, if the confidence threshold is higher than 0.33, the model will reject the sample, failing to provide valuable information to alert healthcare professionals about the patient’s condition, even though the model has enough information to conclude that the tumor is malignant with high certainty. This could potentially result in delayed diagnosis and treatment, posing a significant risk to the patient’s life. However, a hierarchically-aware selective model with an identical confidence threshold would classify the tumor as malignant with 99% confidence. Although this prediction is less specific, it remains highly valuable as it can promptly notify healthcare professionals, leading to early diagnosis and life-saving treatment for the patient.
This motivates us to propose hierarchical selective classification (HSC), an extension of selective classification to a setting where the classes are organized in a hierarchical structure. Such hierarchies are typically represented by a tree-like structure, where each node represents a class, and the edges reflect a semantic relationship between the classes, most commonly an ’is-a’ relationship. Datasets with an existing hierarchy are fairly common, as there are many well-established predefined hierarchies available, such as WordNet [33] and taxonomic data for plants and animals. Consequently, popular datasets like ImageNet [11], which is widely used in computer vision tasks, and iNaturalist [24], have been built upon these hierarchical structures, providing ready-to-use trees. The ImageNet dataset, for example, is organized according to the WordNet hierarchy. A visualization of a small portion of the ImageNet hierarchy is shown in Figure 1. In cases where a hierarchical tree structure is not provided with the dataset, it can be automatically generated using methods such as leveraging LLMs [7, 17, 28, 57].
@@
(4) We conduct a comprehensive empirical study evaluating HSC on more than 1,000 ImageNet classifiers. We present numerous previously unknown observations, most notably that training approaches such as pretraining on larger datasets, contrastive language-image pretraining (CLIP) [35], and knowledge distillation significantly boost selective hierarchical performance. Furthermore, we define hierarchical calibration and discover that HSC consistently improves confidence calibration.
@@ 2 Problem Setup
Selective Classification: Let $\mathcal{X}$ be the input space, and $\mathcal{Y}$ be the label space. Samples are drawn from an unknown joint distribution $P(\mathcal{X}\times\mathcal{Y})$ over $\mathcal{X}\times\mathcal{Y}$. A classifier $f$ is a prediction function $f:\mathcal{X}\rightarrow\mathcal{Y}$, and $\hat{y}=f(x)$ is the model’s prediction for $x$. The true risk of a classifier $f$ w.r.t. $P$ is defined as: $R(f|P)=\mathbb{E}_{P}[\ell(f(X),Y)]$, where $\ell:\mathcal{Y}\times\mathcal{Y}\rightarrow\mathbb{R}^{+}$ is a given loss function, for instance, the 0/1 loss. Given a set of labeled samples $S_{m}=\{(x_{i},y_{i})\}_{i=1}^{m}$, the empirical risk of $f$ is: $\hat{R}(f|S_{m})=\frac{1}{m}\sum_{i=1}^{m}\ell(f(x_{i}),y_{i}).$ Following the notation of [21], we use a confidence score function $\kappa(x,\hat{y}|f)$ to quantify prediction confidence. We require $\kappa$ to induce a partial order over instances in $\mathcal{X}$. In this work, we focus on the most common and well-known $\kappa$ function, softmax response [10, 40]. For a classifier $f$ with softmax as its last layer: $\kappa(x,\hat{y}|f)=f_{\hat{y}}(x)$. Softmax response has been shown to be a reliable confidence score in the context of selective classification [18, 19] as well as in hierarchical classification [46], consistently achieving solid performance.
@@
Hierarchical Classification: Following the notations of [12] and [46], a hierarchy $H=(V,E)$ is defined by a tree, with nodes $V$ and edges $E$, the root of the tree is denoted $r\in V$. Each node $v\in V$ represents a semantic class: the leaf nodes $\mathcal{L}\subseteq V$ are mutually exclusive ground-truth classes, and the internal nodes are unions of leaf nodes determined by the hierarchy. The root represents the semantic class containing all other objects. A sample $x$ that belongs to class $y$ also belongs to the ancestors of $y$. Each node has exactly one parent and one unique path to it from the root node. The set of leaf descendants of node $v$ is denoted by $\mathcal{L}(v)$, and the set of ancestors of node $v$, including $v$ itself, is denoted by $\mathcal{A}(v)$. A hierarchical classifier $f:\mathcal{X}\rightarrow V$ labels a sample $x\in\mathcal{X}$ as a node $v\in V$, at any level of the hierarchy, as opposed to a flat classifier, that only predicts leaves.
@@
Hierarchical Selective Classification: Selective classification offers a binary choice: either to predict or completely reject a sample. We propose hierarchical selective classification (HSC), a hierarchical extension of selective classification, which allows the model to retreat to less specific nodes in the hierarchy in case of uncertainty. Instead of rejecting the whole prediction, the model can now partially reject a sample, and the degree of the rejection is determined by the model’s confidence. For example, if the model is uncertain about the specific dog breed of an image of a Labrador but is confident enough to determine that it is a dog, the safest choice by a classic selective framework would be to reject it. In our setting, however, the model can still provide useful information that the object in the image is a dog (see Figure 1).
@@
To our knowledge, controlling the trade-off between accuracy and specificity was only previously explored by [12], who proposed the Dual Accuracy Reward Trade-off Search (DARTS) algorithm, which aims to obtain the most specific classifier for a user-specified accuracy constraint. They used information gain and hierarchical accuracy as two competing factors, integrating them into a generalized Lagrange function. Our approach differs from theirs in that we guarantee control over the full trade-off, while some coverages cannot be achieved by DARTS.
As mentioned in section 2, it is considered correct to label an image as either its ground truth leaf node or any of its ancestors. Thus, we employ a natural extension of the 0/1 loss to define the true risk of a hierarchical classifier $f^{H}$ with regard to $P$: $R^{H}(f^{H}|P)=\mathds{E}_{P}[f^{H}(X)\notin\mathcal{A}(Y)],$ and the empirical risk over a labeled set of samples $S_{m}$: $\hat{R^{H}}(f^{H}|S_{m})=\frac{1}{m}\sum_{i=1}^{m}\mathds{1}[f^{H}(x_{i})\notin\mathcal{A}(y_{i})].$ An additional risk penalizing hierarchical mistake severity is discussed in Appendix A.
To ensure that specific labels are preferred, it is necessary to consider the specificity of predictions. Hierarchical selective coverage, which we propose as a hierarchical extension of selective coverage, measures the amount of information present in the model’s predictions. A natural quantity for that is entropy, which measures the uncertainty associated with the classes beneath a given node. Assuming a uniform prior on the leaf nodes, the entropy of a node $v\in V$ is $H(v)=-\sum_{{v^{\prime}\in\mathcal{L}(v)}}\frac{1}{|\mathcal{L}(v)|}\log(\frac{1}{|\mathcal{L}(v)|})=\log(|\mathcal{L}(v)|).$ At the root, the entropy reaches its maximum value, $H(r)=\log(|\mathcal{L}|)$, while at the leaves the entropy is minimized, with $H(y)=0$ for any leaf node $y\in\mathcal{L}$. This allows us to define coverage for a single node $v$, regardless of $g_{\theta}^{H}$. We define hierarchical coverage (from now on referred to as coverage) as the entropy of $v$ relative to the entropy of the root node: ${\phi}^{H}(v)=1-\frac{H(v)}{H(r)}.$ The root node has zero coverage, as it does not contain any information. The coverage gradually increases until it reaches 1 at the leaves. We can also define true coverage for a hierarchical selective model: $\phi^{H}(f^{H})=\mathds{E}_{P}[\phi(f^{H}(X)].$ The empirical coverage of a classifier over a labeled set of samples $S_{m}$ is defined as the mean coverage over its predictions: $\hat{\phi^{H}}(f^{H}|S_{m})=\frac{1}{m}\sum_{i=1}^{m}\phi(f^{H}(x_{i})).$ For a hierarchy comprised of leaves and a root node, hierarchical coverage reduces to selective coverage, where classifying a sample as the root corresponds with rejection.
@@
Hierarchical Calibration: Confidence calibration refers to the task of predicting probability estimates that accurately reflect the true likelihood of correctness. As reducing hierarchical coverage may improve accuracy, it may also improve calibration over the samples the model has not abstained from. To capture this relationship, we introduce the concept of Calibration-Coverage Curves (CC curves), which, analogous to RC curves, plot the Expected Calibration Error (ECE) [34] as a function of coverage. see Figure 5 for an example.
@@ 3 Hierarchical Selective Inference Rules
To define a hierarchical selective model $f^{H}=(f,g_{\theta}^{H})$ given a base classifier $f$, an explicit definition of $g_{\theta}^{H}$ is required. $g_{\theta}^{H}$ determines the internal node in the hierarchy $f^{H}$ retreats to when the leaf prediction of $f$ is uncertain. $g_{\theta}^{H}$ requires obtaining confidence for internal nodes in the hierarchy. Since most modern classification models only assign probabilities to leaf nodes, we follow [12] by setting the probability of an internal node to be the sum of its leaf descendant probabilities: $f^{H}_{v}(x)=\sum_{y\in\mathcal{L}(v)}f_{y}(x).$ Unlike other works that assign the sum of leaf descendant probabilities to internal nodes, our algorithm first calibrates leaf probabilities through temperature scaling. Since the probabilities of internal nodes heavily rely on the leaf probabilities supplied by $f$, calibration is beneficial for obtaining more reliable leaf probabilities, and since the internal nodes probabilities are sums of leaf probabilities, this cumulative effect is even more pronounced. Further, [18] found that applying temperature scaling also improves ranking and selective performance, which is beneficial to our objective. In this section, we introduce several inference rules, along with useful theoretical properties. We propose the Climbing inference rule (Algorithm 1), which starts at the most likely leaf and climbs the path to the root, until reaching an ancestor with confidence above the threshold. A visualization of the Climbing inference rule is shown in Figure 1.
We compare our proposed inference rules to the following baselines: The first is the selective inference rule, which predicts ’root’ for every uncertain sample (this approach is identical to standard “hierarchically-ignorant” selective classification). The second hierarchical baseline, proposed by [46], selects the node with the highest coverage among those with confidence above the threshold. We refer to this rule as “Max-Coverage” (MC), and its algorithm is detailed in Appendix B.
@@
The Climbing inference rule outlined in Algorithm 1 satisfies both monotonicity properties. MC satisfies monotonicity in coverage, but not in correctness. An additional algorithm that does not satisfy any of the properties is discussed in Appendix C.
@@ 4 Optimal Selective Threshold Algorithm
3:Based on $\alpha,n,\delta$, calculate $\epsilon$triangleright See Appendix E
@@
6:     $\theta_{i}\leftarrow(g_{\theta}^{H})^{-1}(x_{i}|\kappa,f)$triangleright See Algorithm 5 in Appendix D
@@
Proof: See Appendix E.
Remark on Theorem 1: our algorithm can provide an even greater degree of flexibility: the user may choose to set the values of any three parameters out of $\alpha,n,\epsilon,\delta$. With the three parameters fixed, we can compute the remaining parameter. The size of the calibration set is of particular interest. Although increasing $n$ yields more stable results, our guarantees hold for calibration sets of any size. Even for a small calibration set, our algorithm provides users with the flexibility to set the other parameters according to their requirements and constraints. For example: a user with a low budget for a calibration set who may be, perhaps, more interested in controlling $\delta$ can still achieve a reasonable constraint by relaxing $\epsilon$ instead of increasing the calibration set size. See Appendix E for details.
Table 1: Comparison of mean hAURC and hierarchical gain results for the selective, Max-Coverage (MC) and Climbing inference rules applied to 1,115 models trained on ImageNet1k, and 6 models trained on iNat21.
ImageNet-1k (1,115 models) iNat21 (6 models)
@@
5 Experiments
In this section we evaluate the methods introduced in Section 3 and Section 4. The evaluation was performed on 1,115 vision models pretrained on ImageNet1k [11], and 6 models pretrained on iNat-21 [24] (available in timm 0.9.16 [49] and torchvision 0.15.1 [32]). The reported results were obtained on the corresponding validation sets (ImageNet1k and iNat-21). The complete results and source code necessary for reproducing the experiments are provided in the Supplementary Material.
Inference Rules: Table 1 compares the mean results of the Climbing inference rule to both hierarchical and non-hierarchical baselines. The evaluation is based on RC curves generated for 1,115 models pretrained on ImageNet1k and 6 models pretrained on iNat21 with each inference rule applied to their output. These aggregated results show the effect of different hierarchical selective inference rules, independent of the properties or output of a specific model. Compared to the non-hierarchical selective baseline, hierarchical inference has a clear benefit. Allowing models to handle uncertainty by partially rejecting a prediction instead of rejecting it as a whole, proves to be advantageous; the average model is capable of leveraging the hierarchy to predict internal nodes that reduce risk while preserving coverage. Nonetheless, the differences remain stark when comparing the hierarchical inference rules. Climbing, achieving almost 15% hierarchical gain, outperforms MC, with more than double the gain of the latter. These results highlight the fact that the approach taken by inference rules to leverage the hierarchy can have a significant impact. A possible explanation for Climbing’s superior performance could stem from the fact that most models are trained to optimize leaf classification. By starting from the most likely leaf, Climbing utilizes the prior knowledge embedded in models for leaf classification, while MC ignores it.
Optimal Threshold Algorithm: We compare our algorithm to DARTS [12]. We evaluate the performance of both algorithms on 1,115 vision models trained on ImageNet1k, for a set of target accuracies. For each model and target accuracy the algorithm was run 1000 times, each with a randomly drawn calibration set. We also evaluate both algorithms on 6 models trained on the iNat21 dataset. We compare the algorithms based on two metrics: (1) target accuracy error, i.e., the mean distance between the target accuracy and the accuracy measured on the test set; (2) coverage achieved by using the algorithms’ output on the test set. The results presented in Table 2, and Table 5 in Appendix G, show that our algorithm consistently achieves substantially lower target accuracy errors, indicating that our algorithm succeeds in nearing the target accuracy more precisely. This property allows the model to provide a better balance between risk and coverage. Our algorithm is more inclined towards this trade-off, as it almost always achieves higher coverage than DARTS. This is particularly noteworthy when the target accuracy is high: while DARTS loses coverage quickly, our algorithm maintains coverage that is up to twice as high. Importantly, DARTS does not capture the whole risk-coverage trade-off. Specifically, at the extreme point in the trade-off when it aims to maximize specificity, it still falls short of providing full coverage, and its predictions do not reduce to a flat classifier’s predictions, i.e. it may still predict internal nodes.